# Plugging Attention into Power Grids: Towards Transparent Forecasting

### v1

@@ Abstract
Accurate electricity consumption forecasting is crucial for ensuring grid stability and optimizing power generation, particularly in increasingly decentralized and complex systems. While classical approaches such as Generalized Additive Models (GAMs) remain widely used, they often fail to capture the spatial dependencies inherent in energy networks. Graph Neural Networks (GNNs) offer a principled framework to incorporate this structure by directly leveraging graph topologies. In this work, we evaluate a broad set of GNN architectures – including GCN, GraphSAGE, ChebConv, TAG, APPNP, TransformerConv, and Graph Attention Networks (GAT and GATv2) – on two real-world electricity consumption datasets from France and the UK. Our experiments show that while complex architectures like GATv2 and TransformerConv do not consistently outperform their simpler counterparts, models such as GCN and APPNP achieve strong results in low-data or highly disaggregated settings. Nonetheless, the vanilla GAT remains highly competitive across both datasets and offers an additional interpretability layer via attention mechanisms. We perform a temporal analysis of attention weights, revealing evolving patterns of regional interaction linked to seasonal and meteorological variability. These results highlight that, although attention is not universally superior, it provides valuable explanatory power when spatial dependencies are prominent. Finally, we benchmark ensemble-based expert aggregation strategies, showing that uniform or learned combinations can enhance robustness and outperform individual models under data heterogeneity.
@@
Graph Neural Networks Attention Mechanisms Electricity Load Forecasting Interpretable Machine Learning
@@ 1 Introduction
The effective operation of the electrical system relies on maintaining a balance between electricity supply and demand. Since electricity cannot be efficiently stored, its production needs to be constantly adjusted to match consumption. Providing accurate forecasts for short-term electricity demand is therefore crucial for all participants in the energy market. The shift toward a decentralized electricity network introduces new uncertainties, which pose additional challenges for demand forecasting. In addition, the increasing contribution of renewable energy sources, like solar and wind power, brings fluctuations and intermittency to the electricity market. These fluctuations and intermittency occur at various spatial scales due to the presence of wind farms and photovoltaic power plants. The crisis brought by Covid-19, along with the current economic downturn, add further complexity to forecasting due to the non-stationarity in consumption patterns [2]. The availability of new geolocated data and individual electricity consumption data can be exploited by models that are able to take advantage of additional information and help in minimizing forecast uncertainty [30, 37]. Furthermore, recent advancements in adaptive forecasting algorithms have demonstrated improvements in forecasting quality, particularly for aggregate load forecasting [6, 3].
@@
Deep learning methods have shown great flexibility and adaptability to a wide range of problems applied to the energy sector [29]. To outperform standard methods, they usually require a large amount of data along with intricate relationships between the input features and the target. These conditions are met particularly in decentralized grids, where multiple forecasting problems can be factored with the adequate neural network architecture. Recurrent neural networks (RNNs), particularly Long Short-Term Memory (LSTM) networks, have been widely used in electricity forecasting due to their capacity to model temporal dependencies. [33] proposed a Deep LSTM model for short-term household load forecasting, which showed superior performance over traditional statistical methods such as ARIMA. Similarly, [28] developed a sequence-to-sequence LSTM model that captures the sequential nature of load data effectively. GNNs have gained significant traction for time-series forecasting tasks, due to their ability to model complex spatial-temporal dependencies. In traffic forecasting, for example, [16] proposed a spatio-temporal GNN with attention mechanisms to dynamically weigh spatial and temporal features, significantly improving predictive accuracy. Similarly, [9] developed a dynamic spatio-temporal GNN that constructs adaptive graphs to better capture temporal evolution in traffic flow. In the energy domain, [22] introduced a hybrid GNN framework for the Baidu KDD Cup 2022 wind power challenge, demonstrating the effectiveness of GNNs in modeling power-related spatial dependencies. A broader GNN-based approach was proposed by [21], where external contextual information was integrated into the model to enhance load forecasting robustness. Notably, our previous work [8] was among the first to explore GNNs specifically for electricity consumption forecasting. By leveraging graph-based representations of aggregated load data, we showed that GNNs can outperform classical baselines while revealing interpretable spatial structures across regions. Attention mechanisms have become a central component in deep learning, particularly in sequence modeling and structured data processing. Initially introduced in the context of machine translation [4, 35], attention enables models to dynamically focus on the most relevant parts of the input. In the context of graph machine learning, Graph Attention Networks (GAT) [36] extend this idea by assigning learnable weights to neighboring nodes, allowing for adaptive neighborhood aggregation. More recent variants, such as GATv2 [7] and TransformerConv [34], aim to address limitations of the original formulation by improving expressiveness or robustness. Despite their popularity, several works have raised concerns about the interpretability of attention weights [20], highlighting the need for careful evaluation when using them as explanatory tools. In this work, we leverage attention scores not only for improving predictions but also for extracting interpretable spatio-temporal patterns in power networks. These contributions highlight the promise of GNNs in energy forecasting. However, a key challenge remains: designing or inferring graph structures that are tailored to the forecasting task, rather than relying on fixed or physically-defined topologies.
@@
  • We demonstrate that attention coefficients from GATs provide interpretable, time-varying insights into regional interactions. By applying dimensionality reduction techniques (e.g. PCA and UMAP) to sequences of attention matrices, we reveal coherent seasonal and spatial patterns aligned with physical and meteorological drivers.
@@ 2.2 Graph Convolutions
[…]
@@
$\mathbf{h}_{v}^{(\ell+1)}=\bigg{|}\bigg{|}_{k=1}^{K}\sigma\left(\sum_{u\in\mathcal{N}_{v}}\alpha_{uv}^{(k)}\mathbf{W}^{(k)}\mathbf{h}_{v}^{(\ell)}+\textbf{b}\right),$
@@
GATv2 [7] extends GAT by improving the expressiveness of the attention mechanism. It overcomes the static linear projection limitation of the original GAT by allowing the attention score to depend on both the source and target features after the linear transformation. This offers improved flexibility and results in many tasks. The node update remains similar to GAT, with the attention score computed as:
$\alpha_{uv}^{(k)}=\text{softmax}_{u}\left(\text{LeakyReLU}((\mathbf{a}^{(k)})^{\top}\mathbf{W}^{(k)}[\mathbf{h}_{v}\,||\,\mathbf{h}_{u}])\right).$
@@
$\mathbf{h}_{v}^{(\ell+1)}=\sigma\left((1-\alpha)\sum_{u\in\mathcal{N}_{v}}\hat{A}_{uv}\mathbf{h}_{u}^{(\ell)}+\alpha\mathbf{h}_{v}^{(0)}+\mathbf{b}\right),$ (2)
where $\hat{\mathbf{A}}\in\mathbb{R}^{n\times n}$ denotes the symmetrically normalized adjacency matrix with added self-loops, and $\alpha\in[0,1)$ is the teleport (or restart) probability that controls the strength of the personalized component in the propagation. In our application, we omit the MLP transformation and directly set $\mathbf{H}^{(0)}=\mathbf{X}$ [^4] Pytorch Geometric/APPNP, i.e. the initial node features, which simplifies the architecture while preserving the core propagation mechanism of APPNP.
3 Developing Robust Models with GNNs
This section presents our methodology for building robust GNNs for load forecasting: from the inference of suitable graphs, to the explanation of the forecasts obtained by the models.
3.1 Inferring Graphs from Data
Geographical Data.
@@ 3.1 Inferring Graphs from Data
In practice, $\lambda$ is picked such that the graph induced by $\boldsymbol{W}_{\lambda}$ is minimally connected, and $\sigma_{b}$ is set equal to the median pairwise geodesic distance. Other heuristics can be used to calculate the bandwidth of the Gaussian kernel, but in practice this one gives good results and is robust to outliers [27, 14]. Note that this approach does not take geographic distances into account: we are using a simple approach here, but we could consider reducing the weights of regions separated by physical obstacles (sea, mountains, etc.).
Electricity and Weather Data.
Another way of capturing relationships between nodes is to exploit their spatio-temporal characteristics, particularly those known to influence electricity consumption, such as temperature, cloud cover, or directly the load signal. In practice, to construct the graph, we aggregate the normalized features by region, resulting in $n$ matrices of dimension $d\times T$, where $n$ is the number of nodes, $d$ the number of features, and $T$ the number of time steps. Rather than applying a dimensionality reduction technique, such as the singular value decomposition (SVD) used in [8]), we retain only the most informative signal –the electricity load– to represent each region. This yields a global matrix of dimension $n\times T$ that captures the temporal dynamics of consumption across nodes and serves as a basis for subsequent graph construction algorithms.
In this study, we evaluated three approaches for constructing the weight matrix based on the temporal similarity of regional load signals. The first method computes the Dynamic Time Warping (DTW) distance between regions using the FastDTW algorithm [32], which provides an efficient approximation with linear time and memory complexity. The second method relies on the empirical Pearson correlation matrix computed over the same time-series, where higher correlations suggest stronger connectivity. Finally, the third method uses the inverse of the covariance matrix –known as the precision matrix– which captures partial correlations and is particularly useful for identifying conditional dependencies between nodes. Each of these approaches yields a weighted adjacency matrix that can be used to define the graph structure for subsequent learning.
@@ 4.1 Datasets
We evaluate GNN models on regional electricity consumption data provided by RTE (Réseau de Transport d’Électricité), the French electricity transmission system operator. The dataset spans from January 2015 to December 2019 and includes various features relevant to electricity demand and meteorological conditions across French administrative regions. Calendar-related features include a trend index, multiple date encodings, month and year indicators, time-of-day and day-of-year information, week numbers, day type (weekday, weekend, public holiday), specific holiday periods (e.g. Christmas, summer, national holidays), and a daylight saving time indicator. The target variable is the national electricity load (in megawatts). We also include lagged consumption features, as detailed in Sec. 4.2.
@@
In order to analyze the behavior and interpretability of attention weights in GATs, we generate synthetic datasets specifically designed for structural analysis rather than forecasting performance. Each dataset consists of $N$ time-series, one per node in the graph, defined as the sum of a smooth trend, a periodic component, and additive high-frequency noise. The noise is intentionally strong to prevent models from simply fitting the individual signals, thereby encouraging the discovery of meaningful cross-node dependencies. To simulate interactions between nodes, we transform the original input signals $\mathbf{X}\in\mathbb{R}^{T\times N}$ by $\mathbf{A}\mathbf{X}^{\top}$, where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a coupling matrix. The supervised learning objective is to reconstruct these signals from the initial inputs. This ensures that the predictive performance essentially hinges on leveraging the coupling structure implicitly embedded in the data. We construct three variations of synthetic data:
@@
  • In the explicit switching setting, two distinct coupling matrices $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ are used at different time steps, with an binary input indicating the active matrix ($1$ for $\mathbf{A}_{1}$, 00 for $\mathbf{A}_{2}$). The switching signal is generated as a binary random vector of length $T$, where each value is independently drawn from a uniform distribution over $\{0,1\}$. This results in a stochastic alternation between $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ at each time step (e.g. 0101110101001).
@@ 4.2 Experimental Settings
Let $\mathbf{X}\in\mathbb{R}^{n\times d\times T}$ denote the input tensor of historical node features, and let $\mathbf{y}\in\mathbb{R}^{n\times T}$ represent the corresponding target values, where $n\in\mathbb{N}$ is the number of nodes in the graph, $d\in\mathbb{N}$ is the number of features per node, and $T\in\mathbb{N}$ is the total number of time steps. Prior to training, each input channel is normalized to the range $[0,1]$, and the resulting scaled version is denoted by ~$\tilde{}\mathbf{X}$. We formulate the forecasting task as a sequence-to-sequence regression problem, where the objective is to predict the next $48$ time steps based on the past $48$ observations. More formally, given the input sequence ~$\tilde{}\mathbf{X}_{t-48:t-1}:=\{\tilde{}\mathbf{X}_{\tau}\}_{t-48\leq\tau\leq t-1}$, the model outputs predictions ^$\hat{}\mathbf{y}_{t:t+47}:=\left\{\hat{}\mathbf{y}_{\tau}\right\}_{t\leq\tau\leq t+47}$ via a GNN $\Phi_{\theta}$ defined over the graph $\mathcal{G}$, i.e. ^$\hat{}\mathbf{y}_{t:t+47}=\Phi_{\theta}(\mathcal{G},\tilde{}\mathbf{X}_{t-48:t-1})$. The training objective consists in identifying the parameter set $\theta^{*}$ that minimizes the mean squared error (MSE) loss function $\ell$, formally: $\theta^{*}\in\arg\min_{\theta}~{}\ell(\Phi_{\theta}(\mathcal{G},\tilde{\mathbf{X}}),~{}\mathbf{y})=\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{n}\left(y_{i,t}-\hat{y}_{i,t}\right)^{2}$. The parameters are optimized using mini-batch stochastic gradient descent with the Adam optimizer [23], which provides adaptive learning rates and momentum-based updates for improved convergence.
@@
We evaluated three baseline forecasting strategies to predict the aggregated electricity demand at multiple consumption nodes. The first approach trains a separate feedforward neural network for each node to forecast the next 48 half-hourly values (i.e. the next day), using all available information up to time $t-1$. The second baseline is a classical statistical method: we fit 48 independent SARIMA models per node – one for each half-hour slot – using only historical values (lags), without any exogenous covariates. This approach follows standard practices in time-series modeling that emphasize seasonality and decomposition [5]. The third baseline is a simple yet competitive persistence method, where the forecast for each time slot is directly copied from the same half-hour either on the previous day ($D-1$) or the previous week ($D-7$), a common benchmark in load forecasting studies [18]. These baselines provide a diverse set of references spanning machine learning, statistical, and naive forecasting paradigms.
@@
We evaluate forecasting performance using two standard loss functions: Mean Absolute Percentage Error (MAPE) and Root Mean Square Error (RMSE). MAPE offers an interpretable percentage-based error metric, useful for comparing across scales:
$\mathsf{MAPE}(\mathbf{y},\hat{}\mathbf{y})=\dfrac{1}{T}\sum_{t=1}^{T}\left|{\textstyle\dfrac{\sum_{i=1}^{N}y_{i,t}-\hat{y}_{i,t}}{\sum_{i=1}^{N}y_{i,t}}}\right|\qquad\qquad\mathsf{RMSE}(\mathbf{y},\hat{}\mathbf{y})=\sqrt{\dfrac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{N}(y_{i,t}-\hat{y}_{i,t})^{2}}$ (4)
@@
To ensure fair and performant comparisons across the studied graph-based architectures, we optimized the model hyperparameters using the Optuna library [1]. The search space included core architectural parameters common to all models, such as the number of layers, the number of hidden channels per layer, the batch size, and the learning rate. Each of these plays a distinct role: n_layers controls the receptive field and model depth, hidden_channels regulates the model capacity, batch_size affects the training stability and generalization, and lr determines the speed and stability of convergence. We further included model-specific parameters to adapt the search space to each architecture. For spectral and spatial methods, such as ChebConv and TAGConv, we optimized the neighborhood order (e.g. Chebyshev polynomial degree or number of hops), which modulates the locality of the convolutional filters. For attention-based models such as GAT and TransformerConv, we additionally tuned the number of attention heads, which controls the diversity of subspaces in which attention scores are computed. For APPNP, we included the number of propagation steps $K$, which determines the extent of information diffusion, and the teleport (or restart) probability $\alpha$, which balances the trade-off between local propagation and retention of the original signal. All experiments were performed using early stopping based on validation loss, and each configuration was evaluated through multiple runs. This principled tuning procedure enabled the selection of performant and robust architectures across both datasets. It is also important to note that the graph construction methods differ across datasets: in the French dataset, regional aggregation allows for the meaningful use of spatial graphs informed by physical proximity or meteorological similarity. In contrast, for the UK dataset – which involves highly disaggregated residential-level signals – using a fixed spatial graph can be counterproductive, as the spatial structure is weakly informative or even noisy. In this case, we relied on data-driven graph construction strategies better suited to fine-grained forecasting. The results of the optimization are reported in Table 1.

### v2

@@ Abstract
Reliable prediction of electricity demand plays a key role in safeguarding grid stability and guiding generation decisions, a need that grows with the decentralization and complexity of modern systems. While classical approaches such as Generalized Additive Models (GAMs) remain widely used, they often fail to capture the spatial dependencies inherent in energy networks. Graph Neural Networks (GNNs) offer a principled framework to incorporate this structure by directly leveraging graph topologies. In this work, we evaluate a broad set of GNN architectures —including GCN, GraphSAGE, ChebConv, TAG, APPNP, TransformerConv, and Graph Attention Networks (GAT and GATv2)— on two real-world electricity consumption datasets from France and the UK. Our results show that simpler models such as GCN, SAGE, or APPNP often outperform more complex alternatives in low-data regimes, while GAT ranks among the strongest architectures in our benchmarks, combining high accuracy with valuable interpretability. We perform a temporal analysis of attention weights, revealing evolving patterns of regional interaction linked to seasonal and meteorological variability. These results highlight that, although attention is not universally superior, it provides valuable explanatory power when spatial dependencies are prominent. Additionally, we demonstrate that ensemble-based expert aggregation strategies, particularly bottom-up combinations, significantly improve robustness and yield state-of-the-art performance across both datasets. These findings highlight the dual promise of GNNs for accurate and interpretable forecasting, and suggest that architectural simplicity coupled with ensemble methods can provide a practical path forward for transparent energy analytics.
@@
Graph Neural Networks Attention Mechanisms Load Forecasting Expert Aggregation Interpretable Machine Learning
@@ 1 Introduction
The stability of power systems depends critically on maintaining an equilibrium between generation and demand, a task complicated by the fact that electricity cannot be stored efficiently and must therefore be produced in real time. While ongoing research explores ways to enhance the flexibility of various generation technologies, including nuclear energy [5], accurate short-term demand forecasts remain indispensable for market participants and system operators. The transition toward decentralized grid architectures introduces novel sources of uncertainty, adding complexity to forecasting tasks. At the same time, the growing share of intermittent renewable resources, such as wind and solar power, amplifies variability at multiple spatial scales due to the heterogeneous distribution of production sites. External shocks, notably the Covid-19 crisis and subsequent economic downturn, have further disrupted traditional consumption patterns, underscoring the importance of methods capable of handling non-stationary behaviors [2]. Fortunately, the increasing availability of geolocated and disaggregated consumption data provides new opportunities for models that can leverage such information to reduce forecast uncertainty [31, 40]. Parallel advances in adaptive forecasting techniques have also shown promise, demonstrating measurable improvements in predictive accuracy for aggregate electricity demand [7, 3].
@@
Deep learning approaches have proven remarkably versatile in energy-related applications, showing strong capacity to adapt to diverse forecasting problems [29]. Achieving performance gains over traditional statistical models typically requires access to large datasets as well as the ability to capture complex relationships between predictors and targets. Such conditions are particularly present in decentralized power systems, where numerous forecasting tasks can be decomposed and modeled through suitable neural architectures. Among these, recurrent neural networks (RNNs), and especially Long Short-Term Memory (LSTM) networks, have been extensively applied to electricity demand. For instance, [34] introduced a deep LSTM framework for short-term household load forecasting, demonstrating substantial improvements compared with ARIMA models, while [28] proposed a sequence-to-sequence variant that successfully exploits the sequential structure of load time series. GNNs are increasingly recognized as effective architectures for time-series forecasting in energy systems, thanks to their ability to jointly represent spatial and temporal dependencies. Recent advances illustrate their versatility across different contexts. For instance, [36] proposed GLFN-TC, a short-term forecasting model that combines GNN layers with dilated temporal convolutions and residual connections, enabling automatic graph construction and improved predictive performance. In a multi-building setting, [30] designed a temporal GNN incorporating meteorological variables, which achieved competitive accuracy (MAPE of 3.2% at 15 minutes and 4.2% at 24 hours) while remaining robust to missing data. At a larger scale, PowerGNN [37] explicitly leverages network topology by combining GraphSAGE with GRUs, outperforming classical baselines in renewable-intensive scenarios. Beyond these domain-specific developments, earlier contributions in spatio-temporal learning laid the groundwork for applying GNNs to forecasting tasks. In traffic modeling, [16] introduced an attention-based architecture capable of dynamically weighting spatial and temporal information, while [10] proposed a model where the graph structure itself evolves over time to better capture dynamic dependencies. These ideas have since been transferred to the energy domain. For example, [23] developed a hybrid GNN for the Baidu KDD Cup 2022 wind power challenge, demonstrating the capacity of graph-based approaches to capture intricate inter-node relations in large-scale systems. More recently, [22] proposed an extended framework integrating contextual information, further enhancing the robustness and interpretability of load forecasting models. In earlier work, we investigated the use of GNNs for electricity consumption forecasting [9]. That study suggested that graph-based representations of aggregated load data can provide improvements over classical baselines, while also offering insights into spatial structures across regions. Attention mechanisms have become a central component in deep learning, particularly in sequence modeling and structured data processing. Initially introduced in the context of machine translation [4, 38], attention enables models to dynamically focus on the most relevant parts of the input. In the context of graph machine learning, Graph Attention Networks (GAT) [39] extend this idea by assigning learnable weights to neighboring nodes, allowing for adaptive neighborhood aggregation. More recent variants, such as GATv2 [8] and TransformerConv [35], aim to address limitations of the original formulation by improving expressiveness or robustness. Despite their popularity, several works have raised concerns about the interpretability of attention weights [21], highlighting the need for careful evaluation when using them as explanatory tools. In this work, we leverage attention scores not only for improving predictions but also for extracting interpretable spatio-temporal patterns in power networks. These contributions highlight the promise of GNNs in energy forecasting. However, a key challenge remains: designing or inferring graph structures that are tailored to the forecasting task, rather than relying on fixed or physically-defined topologies.
@@
  • We observe that attention coefficients from GATs vary across time and exhibit structure when projected into low-dimensional spaces using PCA or UMAP. These projections reveal seasonal clusters that align with broad weather regimes (e.g. winter, summer). While these patterns do not constitute strict causal explanations, they offer a first step toward interpretable representations of temporal structure in graph-based forecasts and may support downstream tasks such as regime change detection or model analysis.
@@ 2.2 Graph Convolutions
where $\left[\cdot||\cdot\right]$ is a concatenation operator.
@@
$\mathbf{h}_{v}^{(\ell+1)}=\bigg{|}\bigg{|}_{k=1}^{K}\sigma\left(\sum_{u\in\mathcal{N}_{v}}\alpha_{uv}^{(k)}\mathbf{W}^{(k)}\mathbf{h}_{u}^{(\ell)}+\textbf{b}\right),$
@@
GATv2 [8] extends GAT by improving the expressiveness of the attention mechanism. It overcomes the static linear projection limitation of the original GAT by allowing the attention score to depend on both the source and target features after the linear transformation. This offers improved flexibility and results in many tasks. The node update remains similar to GAT, with the attention score computed as $\alpha_{uv}^{(k)}=\text{softmax}_{u}\left(\text{LeakyReLU}((\mathbf{a}^{(k)})^{\top}\mathbf{W}^{(k)}[\mathbf{h}_{v}\,||\,\mathbf{h}_{u}])\right).$
@@
$\mathbf{h}_{v}^{(\ell+1)}=\sigma\left((1-\alpha)\sum_{u\in\mathcal{N}_{v}}\hat{A}_{uv}\mathbf{h}_{u}^{(\ell)}+\alpha\mathbf{h}_{v}^{(0)}+\mathbf{b}\right),$
where $\hat{\mathbf{A}}\in\mathbb{R}^{n\times n}$ denotes the symmetrically normalized adjacency matrix with added self-loops, and $\alpha\in[0,1)$ is the teleport (or restart) probability that controls the strength of the personalized component in the propagation. In our application, we omit the MLP transformation and directly set $\mathbf{H}^{(0)}=\mathbf{X}$, i.e. the initial node features [^4] Pytorch Geometric/APPNP, which simplifies the architecture while preserving the core propagation mechanism of APPNP.
3 Graph Construction and Model Explainability
In this section, we outline our approach to building GNN-based forecasting models, emphasizing two key aspects: the construction of graph structures from different data sources and the interpretation of the resulting forecasts through explainability techniques.
3.1 Graph Construction from Data
Geographical similarity.
@@ 3.1 Graph Construction from Data
In practice, $\lambda$ is picked such that the graph induced by $\boldsymbol{W}_{\lambda}$ is minimally connected, and $\sigma_{b}$ is set equal to the median pairwise geodesic distance.
Load and meteorological signals.
Rather than applying a dimensionality reduction technique, such as the singular value decomposition (SVD) used in [9], we retain only the most informative signal —the electricity load— to represent each region. This yields a global matrix of dimension $n\times T$ that captures the temporal dynamics of consumption across nodes and serves as a basis for subsequent graph construction algorithms. In this study, we evaluated three approaches for constructing the weight matrix based on the temporal similarity of regional load signals. The first method computes the Dynamic Time Warping (DTW) distance between regions using the FastDTW algorithm [33], which provides an efficient approximation with linear time and memory complexity. The second method relies on the empirical Pearson correlation matrix computed over the same time-series, where higher correlations suggest stronger connectivity. Finally, the third method uses the inverse of the covariance matrix —known as the precision matrix— which captures partial correlations and is particularly useful for identifying conditional dependencies between nodes. Each of these approaches yields a weighted adjacency matrix that can be used to define the graph structure for subsequent learning.
@@ 4.1 Datasets
We evaluate GNN models on regional electricity consumption data provided by RTE (Réseau de Transport d’Électricité), the French electricity transmission system operator. The dataset spans from January 2015 to December 2019 and includes various features relevant to electricity demand and meteorological conditions across French administrative regions. Calendar-related features include a trend index, multiple date encodings, month and year indicators, time-of-day and day-of-year information, week numbers, day type (weekday, weekend, public holiday), specific holiday periods (e.g. Christmas, summer, national holidays), and a daylight saving time indicator. Meteorological features (e.g. temperature, wind, solar irradiance) are averaged across the 32 national weather stations and spatially aggregated by administrative region, providing region-level summaries of environmental conditions. The target variable is the national electricity load (in megawatts). We also include lagged consumption features, as detailed in Sec. 4.2.
@@
In order to analyze the behavior and interpretability of attention weights in GATs, we generate synthetic datasets specifically designed for structural analysis rather than forecasting performance. Each dataset consists of $N$ time-series, one per node in the graph, defined as the sum of a smooth trend, a periodic component, and additive high-frequency noise. The noise is intentionally strong to prevent models from simply fitting the individual signals, thereby encouraging the discovery of meaningful cross-node dependencies. To simulate interactions between nodes, we transform the original input signals $\mathbf{X}\in\mathbb{R}^{N\times T}$ by $\mathbf{A}\mathbf{X}$, where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a coupling matrix. The supervised learning objective is to reconstruct these signals from the initial inputs. This ensures that the predictive performance essentially hinges on leveraging the coupling structure implicitly embedded in the data. We construct three variations of synthetic data:
@@
  • In the explicit switching setting, two distinct coupling matrices $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ are used at different time steps, with an binary input indicating the active matrix ($1$ for $\mathbf{A}_{1}$, 0 for $\mathbf{A}_{2}$). The switching signal is generated as a binary random vector of length $T$, where each value is independently drawn from a uniform distribution over $\{0,1\}$. This results in a stochastic alternation between $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ at each time step (e.g. 0101110101001).
@@ 4.2 Experimental Settings
Let $\tilde{\mathbf{X}}\in\mathbb{R}^{n\times d\times T}$ denote the tensor of normalized historical node features, where $n$ is the number of nodes, $d$ the number of features per node, and $T$ the number of time steps. The target values are represented by $\mathbf{y}\in\mathbb{R}^{n\times T}$. The forecasting problem is cast as a sequence-to-sequence regression task with an input window of length $w=48$ and an output horizon of length $h=48$. At each time step $t$, the GNN $\Phi_{\boldsymbol{\theta}}$ defined on the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ maps the current covariates $\tilde{\mathbf{X}}_{t}$ together with the recent target history $\mathbf{y}_{t-w:t-1}$ to predictions $\hat{\mathbf{y}}_{t:t+h-1}=\Phi_{\boldsymbol{\theta}}(\mathcal{G},\tilde{\mathbf{X}}_{t}\cup\mathbf{y}_{t-w:t-1})$. Model parameters are trained by minimizing the mean squared error (MSE) between predicted and ground-truth sequences, i.e. $\boldsymbol{\theta}^{*}\in\arg\max_{\boldsymbol{\theta}}\;\sum_{t\in\mathcal{T}}\lVert\Phi_{\boldsymbol{\theta}}(\mathcal{G},~\tilde{\mathbf{X}}_{t}\cup\mathbf{y}_{t-1:t-w})-\mathbf{y}_{t:t+h-1}\rVert^{2}$. The parameters are optimized using mini-batch stochastic gradient descent with the Adam optimizer [24], which provides adaptive learning rates and momentum-based updates for improved convergence. We adopt distinct data splits for the two datasets: for the French dataset, the training period spans from January 1, 2015 to December 31, 2017, followed by a one-year validation set (2018) and a one-year test set (2019). For the UK dataset, we use a shorter window: 10 days of training (February 13 to February 22, 2024), one day for validation (February 23, 2024), and three days for testing (February 24 to 26, 2024).
@@
We evaluated three baseline forecasting strategies to predict the aggregated electricity demand at multiple consumption nodes. The first approach trains a separate feedforward neural network for each node to forecast the next 48 half-hourly values (i.e. the next day), using all available information up to time t1t-1. The second baseline is a classical statistical method: we fit 48 independent SARIMA models per node— one for each half-hour slot —using only historical values (lags), without any exogenous covariates. This approach follows standard practices in time-series modeling that emphasize seasonality and decomposition [6]. The third baseline is a simple yet competitive persistence method, where the forecast for each time slot is directly copied from the same half-hour either on the previous day (D1D-1) or the previous week (D7D-7), a common benchmark in load forecasting studies [19]. These baselines provide a diverse set of references spanning machine learning, statistical, and naive forecasting paradigms. While GAMs are widely used in electricity forecasting for their ability to model nonlinear effects of multiple covariates [9], we did not include them here due to the limited availability of informative exogenous features in our setting —particularly in the UK dataset, where only calendar features and lagged consumption values are available. In such contexts, the strength of GAMs in capturing smooth covariate effects becomes less relevant, and simpler autoregressive strategies often suffice [17].
@@
We evaluate forecasting performance at the top scale using two standard loss functions: Mean Absolute Percentage Error (MAPE) and Root Mean Square Error (RMSE). MAPE offers an interpretable percentage-based error metric, useful for comparing across scales:
$\mathsf{MAPE}(\mathbf{y},\hat{}\mathbf{y})=\dfrac{1}{T}\sum_{t=1}^{T}\left|{\textstyle\dfrac{\sum_{i=1}^{N}y_{i,t}-\hat{y}_{i,t}}{\sum_{i=1}^{N}y_{i,t}}}\right|\qquad\qquad\mathsf{RMSE}(\mathbf{y},\hat{}\mathbf{y})=\sqrt{\dfrac{1}{T}\sum_{t=1}^{T}\left(\sum_{i=1}^{N}y_{i,t}-\hat{y}_{i,t}\right)^{2}}$ (3)
@@
To ensure fair and performant comparisons across the studied graph-based architectures, we optimized the model hyperparameters using the Optuna library [1]. The search space included core architectural parameters common to all models, such as the number of layers, the number of hidden channels per layer, the batch size, and the learning rate. Each of these plays a distinct role: n_layers controls the receptive field and model depth, hidden_channels regulates the model capacity, batch_size affects the training stability and generalization, and lr determines the speed and stability of convergence. We further included model-specific parameters to adapt the search space to each architecture. For spectral and spatial methods, such as ChebConv and TAGConv, we optimized the neighborhood order (e.g. Chebyshev polynomial degree or number of hops), which modulates the locality of the convolutional filters. For attention-based models such as GAT and TransformerConv, we additionally tuned the number of attention heads, which controls the diversity of subspaces in which attention scores are computed. For APPNP, we included the number of propagation steps $K$, which determines the extent of information diffusion, and the teleport (or restart) probability $\alpha$, which balances the trade-off between local propagation and retention of the original signal. All experiments were performed using early stopping based on validation loss, and each configuration was evaluated through multiple runs. This principled tuning procedure enabled the selection of performant and robust architectures across both datasets. It is also important to note that the graph construction methods differ across datasets: in the French dataset, regional aggregation allows for the meaningful use of spatial graphs informed by physical proximity or meteorological similarity. In contrast, for the UK dataset —which involves highly disaggregated residential-level signals— using a fixed spatial graph can be counterproductive, as the spatial structure is weakly informative or even noisy. In this case, we relied on data-driven graph construction strategies better suited to fine-grained forecasting. The results of the optimization are reported in Table 1.